OpenStudy (anonymous):

4 = 3/k - 2 Help please?

OpenStudy (anonymous):

4=3/k-2 +2 +2 3*6=3/k*3 18=k

OpenStudy (anonymous):

4*(k-2)=3 k-2=3/4 k=3/4+2 = ?